# On the Utility of Equivariance and Symmetry Breaking in Deep Learning Architectures on Point Clouds

### v1

@@ Abstract
This paper explores the key factors that influence the performance of models working with point clouds, across different tasks of varying geometric complexity. In this work, we explore the trade-offs between flexibility and weight-sharing introduced by equivariant layers, assessing when equivariance boosts or detracts from performance. It is often argued that providing more information as input improves a model’s performance. However, if this additional information breaks certain properties, such as $\mathrm{SE}(3)$ equivariance, does it remain beneficial? We identify the key aspects of equivariant and non-equivariant architectures that drive success in different tasks by benchmarking them on segmentation, regression, and generation tasks across multiple datasets with increasing complexity. We observe a positive impact of equivariance, which becomes more pronounced with increasing task complexity, even when strict equivariance is not required.
@@ 1 Introduction
The inductive bias of weight sharing in convolutions, as introduced in LeCun et al. (2010) traditionally refers to applying the same convolution kernel (a linear transformation) across all neighborhoods of an image. To extend this to transformations beyond translations, Cohen & Welling (2016) introduced Group Equivariant CNN (G-CNNs), adding group equivariance properties to encompass group actions and have weight-sharing across group convolution kernels. G-CNN layers are explicitly designed to maintain equivariance under group transformations, allowing the model to handle transformations naturally without needing to learn invariance to changes that preserve object identity. Following this work—in the spirit of ’convolution is all you need’ (Cohen et al., 2019), several works emerged like (Ravanbakhsh et al., 2017; Worrall et al., 2017; Kondor & Trivedi, 2018; Bekkers et al., 2018; Weiler et al., 2018; Cohen et al., 2019; Weiler & Cesa, 2019; Bekkers, 2019; Sosnovik et al., 2019; Finzi et al., 2020). Advancements in geometric deep learning have demonstrated the effectiveness of incorporating geometric structure as an inductive bias, which reduces model complexity while enhancing generalization and performance (Bronstein et al., 2021). Thus, incorporating group structure into neural networks has become a promising area of research.
However, there is a growing debate as to whether group structure is overly restrictive and if similar advantages could be obtained by simply adding more data. To address this question, we thoroughly investigate the impact of equivariant versus non-equivariant layers in various computational tasks. We explore the balance between leveraging group structures that can provide powerful inductive biases and maintaining the inherent model flexibility. We implemented a convolutional architecture in which the linear layers are either classical or group convolution layers, whilst the overall architecture remains otherwise identical for fair comparison. We evaluate this model on tasks of varying complexity and the extent to which equivariance is desirable under the task description. Our goal is to understand how these design choices affect performance, generalization capabilities, and computational efficiency.
The contributions of this paper can be summarized as follows:
@@
  • Empirical study: To test these hypotheses, we conduct experiments on the point cloud datasets Shapenet 3D, QM9, and CMU Motion Capture across different tasks to assess the empirical effects of using equivariant versus non-equivariant layers.
@@ 2 Background
In this section, we begin by explaining how equivariant neural networks differ from non-equivariant ones, focusing on their architectural distinctions and the impact of weight-sharing—or the lack thereof—on data efficiency. We then delve into the specifics of 3D convolutions and separable group convolutions, which inform the design of the architecture presented later.
@@ 2.1 Arguments for and against equivariance
This paper explores the scenarios in which G-CNNs outperform standard CNNs. We hypothesize that equivariant networks offer advantages not only in tasks explicitly requiring equivariance but also in challenging tasks where strict equivariance is not necessary.
@@ 2.2 Introduction to Equivariant Neural Networks
To address the above arguments we first provide a high-level introduction to group equivariant convolutions in this section and provide the technical details in the next Section 2.3.
@@
Equivariant Networks and Symmetry-Breaking Relaxing equivariant constraints to improve generalization has been well studied. Finzi et al. (2021) presents RPP, which allows for relaxing architectural constraints of equivariance by introducing neural network priors that allow for approximate equivariance. Prior works like Wang et al. (2022); van der Ouderaa et al. (2022); Pertigkiozoglou et al. (2024) focus on relaxing equivariant constraints and show the benefit in performance on learning approximate or partial equivariance. To understand the effects of explicit symmetry breaking while learning strict equivariance in tasks with different geometric complexity, our approach consists of breaking a) internal symmetry and b) external symmetry. See A.3 for more details.
@@ 2.3 Equivariant Linear Layers (Convolutions)
The general form of a linear layer for such point cloud feature fields is given by (Bekkers, 2019)
@@
In case one does not want to impose any further constraints on the kernel $k$ but still wants to remain fully $\mathrm{SE}(3)$ equivariant, one has no other option than adding an axis over which to organize convolution responses in terms of rotations (Bekkers, 2019, Theorem 1). One then has to utilize lifting convolutions, followed by group convolutions:
lifting conv: $[Lf](\mathbf{x}_{i},\mathbf{R})$ $=\sum_{j\in\mathcal{N}(i)}k(\mathbf{R}^{T}(\mathbf{x}_{j}-\mathbf{x}_{i}))f(\mathbf{x}_{j})\,,$ (5)
group conv: $[Lf](\mathbf{x}_{i},\mathbf{R})$ $=\sum_{j\in\mathcal{N}(i)}\int_{SO(d)}k(\mathbf{R}^{T}(\mathbf{x}_{j}-\mathbf{x}_{i}),\mathbf{R}^{T}\tilde{\mathbf{R}})f(\mathbf{x}_{j},\tilde{\mathbf{R}}){\rm d}\tilde{\mathbf{R}}\,,$ (6)
with ${\rm d}\tilde{\mathbf{R}}$ denoting the Haar measure over the rotation group $SO(3)$. Note that now the convolution kernel in equation 5 is an unconstrained function over $\mathbb{R}^{3}$, and that of Eq. 6 over $\mathbb{R}^{3}\times\mathrm{SO}(3)$, and that this kernel is rotated for every possible $\mathbf{R}\in\mathrm{SO}(3)$. Both the lifting and group convolution layers generate feature maps $\mathcal{X}\times\mathrm{SO}(3)\rightarrow\mathbb{R}^{C^{\prime}}$ over the joint space of positions $\mathcal{X}\subset\mathbb{R}^{3}$ and rotations $\mathrm{SO}(3)$. In other words, at each point $\mathbf{x}_{i}$ we now have a signal over the rotation group $\mathrm{SO}(3)$ which stores the convolution response for every “pose" $\mathbf{R}$ in $\mathrm{SO}(3)$, and the linear layer is defined using convolution kernels that match feature patterns of relative spatial and rotational poses.
Eqs. 5 and 6 are forms of regular group convolutions Cohen & Welling (2016). Brandstetter et al. show that such layers can also be implemented through tensor field layers (Thomas et al., 2018), which form a popular class of steerable group convolutions that are parametrized by Clebsch-Gordan tensor products and work with vector fields that transform via irreducible representations (Weiler et al., 2021). Following Bekkers et al. (2024), we note however, that tensor-field networks unnecessarily constrain neural network design—as they require specialized activation functions, they require in-depth knowledge of representation theory, and they are computationally demanding due to the use of Glebsch-Gordan tensor products. Our study therefore focuses on regular group convolutions.
@@
with $c,c^{\prime}$ the row and column indices of the "channel mixing" matrix. Then Eq. 6 can be split into three steps that are each efficient to compute: a spatial interaction layer (message passing), a point-wise $\mathrm{SO}(3)$ convolution, and a point-wise linear layer (Knigge et al., 2022; Kuipers & Bekkers, 2023). It would result in the group convolutional counterpart of depth-wise separable convolution (Chollet, 2017), which separates convolution in two steps (spatial mixing and channel mixing).
@@
again, with no channel-mixing taking place. Note that when the sphere $S^{2}$ is discretized with $O$ number of orientations, Eq. 9 is implemented as a point-wise matrix multiplication with a precomputed kernel matrix of size $O\times O\times C$, e.g. via einsum("noc,poc->npc", features, spherical_kernel), with tensors of $f$ being of dimensions $N\times O\times C$. Finally, $L^{(channel)}$ is simply a point-wise linear layer that mixes the channels without any spatial or orientation mixing, i.e., it is distributed over all points ($N$ axis) and orientations ($O$ axis) and thus is efficiently computed.
@@ 3.1 Hypotheses on the impact of Equivariance
However, we also recognize that standard (non-equivariant) neural networks may not need to rely on these properties when a large amount of data is available. In such cases, the first argument no longer applies, and the second argument becomes less significant because a large training set might be representative of the test set—or, in a sense, even include it. [^1] A highly unlikely scenario when dealing with real-world data.. Moreover, data augmentation might be used to promote generalization over symmetries [^2] Moskalev et al. (2023) however empirically showed that at best, an apparent equivariance property is attained, which breaks under distribution shifts.. We thus formulate the following hypothesis, which will be tested with scaling experiments, with data augmentation enabled in both cases.
Figure 1: Progression of tasks based on geometric complexities.
Hypothesis 1 (Scaling laws)
For tasks that require invariance or equivariance, both equivariant and non-equivariant models will converge to similar performance with increasing dataset size.
@@
    In small-scale data regimes equivariant models will outperform non-equivariant models.
We must also consider task complexity, shown in Figure 1. We will focus the study on tasks that require either invariance or equivariance and formulate the following hypothesis to study the role of equivariance in tasks with varying geometric complexity.
Hypothesis 2 (Geometric complexity)
The advantages of equivariant methods over non-equivariant ones become more pronounced in tasks that require a stricter form of equivariance.
@@
    There is a hierarchy of tasks in terms of how geometrically demanding they are, and thus to what extent equivariance is needed.
@@
Hypothesis 3 (Representation capacity)
Under the same representation capacity (e.g. same number of total channels per node), equivariant and non-equivariant models perform on par.
@@
Hypothesis 4 (Kernel constraint)
We expect that unconstrained (translational) models (Eq. 3) outperform (roto-translation) models (Eq. 4) over $\mathbb{R}^{3}$ as the former is less constrained. Similarly, we expect that $\mathrm{SE}(3)$ equivariant group convolution methods (Eq. 7) outperform the constrained $\mathbb{R}^{3}$ convolutional models (Eq. 4), even when matched in representation capacity.
@@
Hypothesis 5 (Symmetry breaking and explicit geometric representations)
In non-equivariant tasks, where symmetry breaking is allowed, models that are provided with explicit geometric information (such as taking coordinates as features) outperform those that do not have this information.
@@ 3.2 Architecture
Based on (Bekkers et al., 2024), we design a model class called Rapidash that is capable of incorporating various forms of equivariances, and thus has the flexibility to test the above hypotheses.
@@
Our architecture is fully convolutional and is based on the ConvNext block from Liu et al. (2022). See 2 for an illustration. The block consists of the following sequence of operations: a depth-wise separable convolution, followed by a LayerNorm, then a point-wise linear layer that increases the channel dimension four-fold, a GELU activation, and a linear back to the original channel dimensions. This output is added via a skip connection. The depth-wise separable convolution is either implemented via Eq. 3 or 4 for the classical (base space $\mathbb{R}^{3}$) architecture, and via Eq. 8 and 9 in the $\mathbb{R}^{3}\times S^{2}$ case. Thus, the only difference to the point-cloud implementation of ConvNext is that in our group convolutional implementation, the convolution is over $\mathbb{R}^{3}\times S^{2}$ instead of just over $\mathbb{R}^{3}$.
@@ 4 Experiments
We present a thorough analysis of the hypotheses mentioned in Section 3 through a series of experiments on various point cloud datasets like Shapenet 3D, QM9, and the CMU human motion prediction dataset, which are tasks with increasing levels of complexity (Fig. 1). For each of the experiments, we have different variations of Rapidash arising from equations Eq. 3 or 4 for $\mathbb{R}^{3}$ and Eq. 8 and 9 in $\mathbb{R}^{3}\times S^{2}$ case, all with different input variations. For each table, the top section shows models with $\mathrm{SE}(3)$ equivariance, and the bottom section shows models with $T_{3}$ equivariance. Symmetry breaking is indicated by: !  ($\mathrm{SE}(3)$), !  ($T_{3}$ or $\mathrm{SO}(3)$), and !  (conditional $\mathrm{SO}(3)$). A indicates an option is used, , indicates it is not used, and "-" indicates the option is not available. ➟  marks the model with maximal equivariance and information access. All experiments use Rapidash under different equivariance constraints. Note, all models have implicit access to positional information via pair-wise geometric attributes. For implementation details see App. A.5. Our code is made available at [see anonymous supplementary .zip file].
4.1 3D point cloud segmentation and generation experiments
In this experiment, we evaluate our models on the ShapeNet 3D dataset (Chang et al., 2015) part segmentation and generation tasks. ShapeNet consists of 16,881 shapes from 16 categories. Each shape is annotated with up to six parts, totaling 50 parts. We use the point sampling of 2,048 points and the train/validation/test split from (Qi et al., 2017). We compare our models to various state-of-the-art methods like PointnXt (Qian et al., 2022), Deltaconv (Wiersma et al., 2022), and GeomGCNN (Srivastava & Sharma, 2021)for the part segmentation task. For the generation task, we compare to the latent diffusion model, LION, (Zeng et al., 2022).
@@
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}80.26_{\pm 0.06}}80.31_{\pm 0.15}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}80.33}_{\pm 0.13}}80.20_{\pm 0.07}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
2 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}83.95_{\pm 0.06}}83.87_{\pm 0.09}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}52.09_{\pm 0.87}}49.63_{\pm 0.87}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
3 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}84.23_{\pm 0.08}}84.01_{\pm 0.06}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}34.15_{\pm 0.05}}32.22_{\pm 0.27}\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.92}_{\pm 0.98}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
4 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}84.75}_{\pm 0.02}}84.48_{\pm 0.16}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}34.07_{\pm 0.43}}32.90_{\pm 0.47}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$

### v2

@@ Abstract
In this work, we explore the trade-offs between flexibility and weight-sharing introduced by equivariant layers, assessing when equivariance boosts or detracts from performance. It is often argued that providing more information as input improves model performance. However, if this additional information breaks certain properties, such as $\mathrm{SE}(3)$ equivariance, does it remain beneficial? We identify the key aspects of equivariant and non-equivariant architectures that drive success in different tasks by benchmarking them on classification, segmentation, regression, and generation tasks across multiple datasets with increasing task complexity. We observe a positive impact of equivariance, which becomes more pronounced with increasing task complexity, even when strict equivariance is not required.
Machine Learning, ICML
@@ 1 Introduction
The inductive bias of weight sharing in convolutions, as introduced in LeCun et al. (2010) traditionally refers to applying the same convolution kernel (a linear transformation) across all neighborhoods of an image. To extend this to transformations beyond translations, Cohen & Welling (2016) introduced Group Equivariant CNN (G-CNNs), adding group equivariance properties to encompass group actions and have weight-sharing across group convolution kernels. G-CNN layers are explicitly designed to maintain equivariance under group transformations, allowing the model to handle transformations naturally without needing to learn invariance to changes that preserve object identity. Following this work—in the spirit of ’convolution is all you need’ (Cohen et al., 2019), several works emerged like (Ravanbakhsh et al., 2017; Worrall et al., 2017; Kondor & Trivedi, 2018; Bekkers et al., 2018; Weiler et al., 2018; Cohen et al., 2019; Weiler & Cesa, 2019; Bekkers, 2019; Sosnovik et al., 2019; Finzi et al., 2020). Advancements in geometric deep learning have demonstrated the effectiveness of incorporating geometric structure as an inductive bias, which reduces model complexity while enhancing generalization and performance (Bronstein et al., 2021).
However, there is an ongoing debate about the necessity of explicit structural priors like group-equivariant architectures for generalization. One perspective emphasizes the power of scaling data and model capacity (Qu & Krishnapriyan, 2024), arguing that strict adherence to group structure may unnecessarily constrain models and limit their scalability—a view supported by recent successes like AlphaFold 3 (Abramson et al., 2024), which deliberately compromises exact equivariance for improved computational efficiency. While proponents argue that sufficient data and appropriate augmentations can naturally induce approximate symmetries, Moskalev et al. (2023) demonstrate that such learned properties are often superficial and inconsistent. This is further supported by theoretical work from Perin & Deny (2024), which proves fundamental limitations on learning equivariance from data alone, particularly when class orbits are insufficiently distinct or aren’t densely packed. Petrache & Trivedi (2023) formalizes these trade-offs between approximate equivariance and generalization, showing that even partial symmetry alignment can improve performance—helping explain why equivariant architectures continue to dominate in symmetry-rich domains like molecular modeling and physical systems, where exact equivariance properties are crucial for reliable predictions.
To address this debate empirically, we systematically investigate the impact of equivariant versus non-equivariant layers across diverse computational tasks. We explore the fundamental trade-off between leveraging group structures as inductive biases and maintaining model flexibility.We implement a convolutional architecture where only the linear layers vary between classical and group convolutions, while keeping all other architectural components identical for controlled comparison. We evaluate this model on tasks of varying complexity and the extent to which equivariance is desirable under the task description. Our goal is to understand how these design choices affect performance, generalization capabilities, and computational efficiency.
The contributions of this paper can be summarized as:
@@
  • Empirical study: To test these hypotheses, we conduct experiments on the point cloud datasets Shapenet 3D Modelnet40, CMU Motion Capture, and QM9 across different tasks.
@@ 2 Background
In this section, we explain how equivariant neural networks differ from non-equivariant ones, discuss their architectural distinctions, and the impact of weight-sharing—or the lack thereof—on data efficiency. Additionally, we discuss 3D convolutions and separable group convolutions in detail to inform the design of the Rapidash architecture used for this study
@@ 2.2 Introduction to Equivariant Neural Networks
In this section, we provide a high-level introduction to group equivariant convolution and provide the technical details in the next Section 2.3.
@@
Equivariant Networks and Symmetry-Breaking
Relaxing equivariant constraints to improve generalization has been well studied. (Finzi et al., 2021) presents RPP, which allows for relaxing architectural constraints of equivariance by introducing neural network priors that allow for approximate equivariance. Prior works like (Wang et al., 2022; van der Ouderaa et al., 2022; Pertigkiozoglou et al., 2024) focus on relaxing equivariant constraints and show the benefit in performance on learning approximate or partial equivariance. These methods break equivariance by loosening equivariance equivariance constraints of layers $L$—a mechanism we refer to as internal symmetry breaking. We observe, however, that equivariance is also broken by miss-specification of inputs (e.g. treating vectors as scalars) whilst keeping the inner workings in tact. We refer to the as external symemtry breaking, see A.3 for more details. Our experiments abblate over the effect of both internal symmetry breaking (miss-specification of equivariance constraints) and external symmetry breaking.
@@ 2.3 Equivariant Linear Layers (Convolutions)
The general form of a linear layer for such point cloud feature fields is given by
@@
In case one does not want to impose any further constraints on the kernel $k$ but still wants to remain fully $\mathrm{SE}(3)$ equivariant, one has no other option than adding an axis over which to organize convolution responses in terms of rotations (Bekkers, 2019, Theorem 1). One then has to utilize lifting convolutions eq.5, followed by group convolutions eq.6:
$[Lf](\mathbf{x}_{i},\mathbf{R})\!$ $=\!\sum_{j\in\mathcal{N}(i)}k(\mathbf{R}^{T}(\mathbf{x}_{j}-\mathbf{x}_{i}))f(\mathbf{x}_{j})\,,$ (5)
$[Lf](\mathbf{x}_{i},\mathbf{R})\!$ $=$ (6)
$\!\sum_{j\in\mathcal{N}(i)}\int\limits_{SO(d)}k(\mathbf{R}^{T}(\mathbf{x}_{j}-\mathbf{x}_{i}),\mathbf{R}^{T}\tilde{\mathbf{R}})f(\mathbf{x}_{j},\tilde{\mathbf{R}}){\rm d}\tilde{\mathbf{R}}\,,$
with ${\rm d}\tilde{\mathbf{R}}$ denoting the Haar measure over the rotation group $SO(3)$. Note that now the convolution kernel in equation 5 is an unconstrained function over $\mathbb{R}^{3}$, and that of Eq. 6 over $\mathbb{R}^{3}\times\mathrm{SO}(3)$, and that this kernel is rotated for every possible $\mathbf{R}\in\mathrm{SO}(3)$. Both the lifting and group convolution layers generate feature maps $\mathcal{X}\times\mathrm{SO}(3)\rightarrow\mathbb{R}^{C^{\prime}}$ over the joint space of positions $\mathcal{X}\subset\mathbb{R}^{3}$ and rotations $\mathrm{SO}(3)$. In other words, at each point $\mathbf{x}_{i}$ we now have a signal over the rotation group $\mathrm{SO}(3)$ which stores the convolution response for every “pose” $\mathbf{R}$ in $\mathrm{SO}(3)$, and the linear layer is defined using convolution kernels that match feature patterns of relative spatial and rotational poses.
Eqs. 5 and 6 are forms of regular group convolutions (Cohen & Welling, 2016). Brandstetter et al. (2022) show that such layers can also be implemented through tensor field layers (Thomas et al., 2018), which form a popular class of steerable group convolutions that are parametrized by Clebsch-Gordan tensor products and work with vector fields that transform via irreducible representations (Weiler et al., 2021). Following (Bekkers et al., 2024), we note however, that tensor-field networks unnecessarily constrain neural network design—as they require specialized activation functions, they require in-depth knowledge of representation theory, and they are computationally demanding due to the use of Glebsch-Gordan tensor products. Our study therefore focuses on regular group convolutions.
@@
with $c,c^{\prime}$ the row and column indices of the ”channel mixing” matrix. Then Eq. 6 can be split into three steps that are each efficient to compute: a spatial interaction layer (message passing), a point-wise $\mathrm{SO}(3)$ convolution, and a point-wise linear layer (Knigge et al., 2022; Kuipers & Bekkers, 2023). It would result in the group convolutional counterpart of depth-wise separable convolution (Chollet, 2017), which separates convolution in two steps (spatial mixing and channel mixing).
Figure 1: Progression of tasks based on geometric complexities.
@@
again, with no channel-mixing taking place. Note that when the sphere $S^{2}$ is discretized with $O$ number of orientations, Eq. 9 is implemented as a point-wise matrix multiplication with a precomputed kernel matrix of size $O\times O\times C$, e.g. via einsum("noc,poc->npc", features,
spherical_kernel), with tensors of $f$ being of dimensions $N\times O\times C$. Finally, $L^{(channel)}$ is simply a point-wise linear layer that mixes the channels without any spatial or orientation mixing, i.e., it is distributed over all points ($N$ axis) and orientations ($O$ axis) and thus is efficiently computed.
@@ 3.1 Hypotheses on the impact of Equivariance
However, we also recognize that standard (non-equivariant) neural networks may not need to rely on these properties when a large amount of data is available. In such cases, the first argument no longer applies, and the second argument becomes less significant because a large training set might be representative of the test set—or, in a sense, even include it. [^1] A highly unlikely scenario when dealing with real-world data.. Moreover, data augmentation might be used to promote generalization over symmetries222(Moskalev et al., 2023) however empirically showed that at best, an apparent equivariance property is attained, which breaks under distribution shifts.. We thus formulate the following hypothesis, which will be tested with scaling experiments with data augmentation enabled in both cases.
Hypothesis 1 [Scaling laws] For tasks that require invariance or equivariance, both equivariant and non-equivariant models will converge to similar performance with increasing dataset size.
@@
    In small-scale data regimes, equivariant models will outperform non-equivariant models.
We must also consider the complexity of the task, shown in Figure 1. We will focus on tasks that require either invariance or equivariance and formulate the following hypothesis to study the role of equivariance in tasks with varying geometric complexity.
Hypothesis 2 [Geometric complexity] The advantages of equivariant methods over non-equivariant ones become more pronounced in tasks that require a stricter form of equivariance.
@@
    There is a hierarchy of tasks regarding how geometrically demanding they are, and thus to what extent equivariance is needed.
@@
Hypothesis 3 [Representation capacity] Under the same representation capacity (e.g. same number of total channels per node), equivariant and non-equivariant models perform on par.
@@
Hypothesis 4 [Kernel constraint] We expect that unconstrained (translational) models (Eq. 3) outperform (roto-translation) models (Eq. 4) over $\mathbb{R}^{3}$ as the former is less constrained. Similarly, we expect that the $\mathrm{SE}(3)$ equivariant group convolution methods (Eq. 7) outperform the constrained $\mathbb{R}^{3}$ convolutional models (Eq. 4), even when matched in representation capacity.
@@
Hypothesis 5 [Symmetry breaking and explicit geometric representations] In non-equivariant tasks, where symmetry breaking is allowed, models that are provided with explicit geometric information (such as taking coordinates as features) outperform those that do not have this information.
@@ 3.2 Architecture
Based on (Bekkers et al., 2024), we design a model class called Rapidash that is capable of incorporating various forms of equivariance, and thus has the flexibility to test the above hypotheses.
@@
Our architecture is fully convolutional and is based on the ConvNext block from (Liu et al., 2022). We illustrate it in Figure 5. The block consists of the following sequence of operations: a depth-wise separable convolution, followed by a LayerNorm, then a point-wise linear layer that increases the channel dimension four-fold, a GELU activation, and a linear back to the original channel dimensions. This output is added via a skip connection. The depth-wise separable convolution is either implemented via Eq. 3 or 4 for the classical (base space $\mathbb{R}^{3}$) architecture, and via Eq. 8 and 9 in the $\mathbb{R}^{3}\times S^{2}$ case. Thus, the only difference to the point-cloud implementation of ConvNext is that in our group convolutional implementation, the convolution is over $\mathbb{R}^{3}\times S^{2}$ instead of just over $\mathbb{R}^{3}$.
@@ 4 Experiments
We present a thorough analysis of the hypotheses mentioned in Section 3 through a series of experiments on various point cloud datasets like Shapenet 3D, ModelNet40 (App. A.5), CMU human motion prediction, and QM9 dataset with tasks of varying levels of complexity (Fig. 1). Here, we emphasize that these experiments aim to demonstrate how a simple group convolutional architecture, combined with symmetry breaking, can achieve comparable performance to state-of-the-art methods, rather than necessarily surpassing them. For each of the experiments, we have different variations of Rapidash arising from equations Eq. 3 or 4 for $\mathbb{R}^{3}$ and Eq. 8 and 9 in $\mathbb{R}^{3}\times S^{2}$ case, all with different input variations. In each table, the top section shows models with $\mathrm{SE}(3)$ equivariance, and the bottom shows models with $T_{3}$ equivariance. Symmetry breaking is indicated by: !  ($\mathrm{SE}(3)$), !  ($T_{3}$ or $\mathrm{SO}(3)$), and !  (conditional $\mathrm{SO}(3)$). A indicates an option is used, , indicates it is not used, and ”-” indicates the option is not available. ➟  marks the model with maximal equivariance and information access. Note, that all models have implicit access to positional information via pair-wise geometric attributes. For implementation details see App. A.6.
Figure 2: Depiction of ShapeNet segmentation for the class airplane. When the sample is rotated, a non-equivariant model mistakes wings for the nose when they get aligned to different axes as they can overfit to global coordinates, while an equivariant model perfectly generalizes over rotations.
Table 1: Ablation studies for ShapeNet3D and CMU Motion Capture dataset. Results on the ShapeNet dataset are reported for part segmentation (mean IoU) on aligned and randomly rotated models, as well as shape generation (1-NNA). For motion prediction MSE(×$\times 10^{-2}$) and rotated MSE (×$\times 10^{-2}$) are reported. Best results are highlighted in blue.
@@
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $79.08_{\pm 0.22}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}79.08}_{\pm 0.33}$ $>100$ $>100$
2 $\mathbb{R}^{3}$ $T_{3}$ $83.54_{\pm 0.04}$ $49.06_{\pm 2.27}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.95}_{\pm 0.11}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}24.62}_{\pm 2.83}$
3 $\mathbb{R}^{3}$ none $83.65_{\pm 0.03}$ $32.14_{\pm 0.17}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}69.19}_{\pm 0.30}$ $6.77_{\pm 0.06}$ $92.10_{\pm 9.41}$
4 $\mathbb{R}^{3}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}84.43}_{\pm 0.11}$ $33.83_{\pm 0.36}$ $6.04_{\pm 0.21}$ $>100$